### v2

# AutoAct: Automatic Agent Learning from Scratch via Self-Planning

Abstract
Language agents have achieved considerable performance on various complex tasks. Despite the incessant exploration in this field, existing language agent systems still struggle with costly, non-reproducible data reliance and face the challenge of compelling a single model for multiple functions. To this end, we introduce AutoAct, an automatic agent learning framework that does not rely on large-scale annotated data and synthetic trajectories from closed-source models (e.g., GPT-4). Given limited data with a tool library, AutoAct first automatically synthesizes planning trajectories without any assistance from humans or strong closed-source models. Then, AutoAct leverages a division-of-labor strategy to automatically differentiate based on the target task information and synthesized trajectories, producing a sub-agent group to complete the task. We conduct comprehensive experiments with different LLMs, which demonstrates that AutoAct yields better or parallel performance compared to various strong baselines. We even notice that AutoAct, when using the Llama-2-13b model, can achieve performance comparable to that of the zero-shot GPT-3.5-Turbo agent [^1] Code will be available at https://github.com/zjunlp/AutoAct..
1 Introduction
Language agents Wang et al. (2023b); Xi et al. (2023); Xie et al. (2023), which leverage the powerful reasoning capabilities Qiao et al. (2023b); Zhang et al. (2023b) of Large Language Models (LLMs) to generate executable actions for observing the external world, have emerged as essential components of AI systems designed to address intricate interactive tasks Torantulino (2023); Osika (2023); Nakajima (2023); Tang et al. (2023). The process of endowing LLMs with such interactive capabilities is referred to as Agent Learning wherein planning plays a pivotal role, which is responsible for decomposing complex tasks Wei et al. (2022); Zhou et al. (2023a); Yao et al. (2023), invoking external tools Shen et al. (2023); Lu et al. (2023); Qin et al. (2023), reflecting on past mistakes Shinn et al. (2023); Madaan et al. (2023), and aggregating information from various sources to reach the final targets. There have been a lot of works Yao et al. (2023); Wang et al. (2023a); Shen et al. (2023); Chen et al. (2023c) that directly prompt closed-source off-the-shelf LLMs to plan on particular tasks. Despite their convenience and flexibility, closed-source LLMs inevitably suffer from unresolved issues, as their accessibility often comes at a steep price and their black-box nature makes the result reproduction difficult. In light of this, some recent endeavors have shifted their focus towards imbuing open-source models with planning capabilities through fine-tuning Chen et al. (2023a); Zeng et al. (2023); Yin et al. (2023).
Figure 1: The basic framework of AutoAct. Armed with just one tool library, the Meta-Agent can automatically differentiate based on the target task information and produce a sub-agent group that can collaborate to complete the task.
Method Data Acquisition Trajectory Acquisition Planning Multi-Agent Fine-Tuning Generality Reflection
  ReAct Yao et al. (2023) User Prompt Iterative
Reflexion Shinn et al. (2023) User Prompt Iterative
Chameleon Lu et al. (2023) User Prompt Global
HuggingGPT Shen et al. (2023) User Prompt Global
AutoGPT Torantulino (2023) User Prompt Iterative
BOLAA Liu et al. (2023) User Prompt Iterative
AgentVerse Chen et al. (2023c) User Prompt Iterative
Agents Zhou et al. (2023c) User Prompt Iterative
AgentTuning Zeng et al. (2023) Benchmark GPT-4 Iterative
FireAct Chen et al. (2023a) Benchmark GPT-4 Iterative
AutoAct (ours) User + Self-Instruct Self-Planning Iterative
Table 1: Comparison of related works. Data and Trajectory Acquisitions refer to the way for obtaining training data and trajectories. Planning represents the way of planning, parted based on whether each step’s action is determined globally or iteratively. Multi-Agent indicates whether the framework contains multi-agent. Fine-Tuning stands for whether the method is a fine-tuning-based agent learning framework. Generality signifies whether the method is applicable to various tasks. Reflection denotes whether the planning process incorporates reflection.
However, despite the existing fine-tuning-based method’s achievements, they are not without their limitations. On the one hand, training open-source models necessitates a substantial amount of annotated task data and still relies on closed-source models to synthesize planning trajectories. However, fulfilling these requirements in many real-world scenarios, such as private personal bot or sensitive company business, often proves to be rocky. On the other hand, from the perspective of agent framework design, fine-tuning-based methods compel one single language agent to learn all planning abilities, placing even greater pressure on them. These contradict the Simon’s principle of bounded rationality Mintrom (2015), which states that “precise social division-of-labor and clear individual tasks can compensate for the limited ability of individuals to process and utilize information”.
To this end, we introduce AutoAct, an automatic agent learning framework, which does not rely on large-scale annotated data and synthetic trajectories from closed-source models while incorporating explicit individual tasks with precise division-of-labor (see Figure 1). Given a limited set of user-provided data examples, AutoAct starts with a Meta-Agent to obtain an augmented database through self-instruct Wang et al. (2023c). Then, armed with a prepared tool library, the Meta-Agent can automatically synthesize planning trajectories without any assistance from humans or strong closed-source models. Finally, we propose the division-of-labor strategy which resembles cell differentiation based on the self-synthesized trajectories (genes), where the Meta-Agent acts as a stem cell Colman (2008) and differentiates into three sub-agents with distinct functions: task decomposition, tool invocation, and self-reflection, respectively. Our differentiation process is essentially a parameter-efficient training process on the self-synthesized trajectories with low consumption resources. We list the differences between AutoAct and prior works in Table 1.
Experiments on complex question-answering tasks with different LLMs demonstrate that the proposed AutoAct yields better or parallel performance compared to various strong baselines. We summarize our main contributions as follows:
  • We propose AutoAct, an automatic agent learning framework that does not rely on large-scale annotated data and synthetic trajectories from closed-source models while adhering to the principle of bounded rationality.
  • We conduct comprehensive experiments with different LLMs, which demonstrates that AutoAct yields better or parallel performance compared to various strong baselines. We even notice that when using the Llama-2-13b model, AutoAct can achieve performance comparable to that of the zero-shot GPT-3.5-Turbo agent OpenAI (2022).
  • Extensive empirical analysis demonstrates the effectiveness of our appropriate division-of-labor strategy and the trajectory quality generated by AutoAct outperforms that of other methods from multiple aspects.
2 AutoAct
Figure 2: The overview of our proposed framework AutoAct. We initiate with self-instruct to extend the task database from scratch. Then self-planning is applied to conduct automatic agent learning, including automatic tool selection, trajectories synthesis, self-differentiation and group planning. Our self-differentiation is a parameter-efficient fine-tuning process to achieve resource-efficient learning.
Overview.
As shown in Figure 2, AutoAct only requires target task information and a language agent (we name it Meta-Agent) to initiate its work. The Meta-Agent first augments the task data from scratch by self-instruct. Furthermore, with a tool library available, the Meta-Agent conducts automatic agent learning by differentiating into sub-agents with distinct functionalities and enabling them to perform group task-specific planning. We name this process as self-planning. Below is a detailed introduction of AutoAct. Note that all symbols used are globally defined.
2.1 Critical Components of AutoAct
Meta-Agent.
The Meta-Agent stands at the central position of our AutoAct framework. It is responsible for all the preparatory work before self-differentiation and serves as the backbone model for the self-differentiation process. Given limited target task information and a pre-prepared tool library, the Meta-Agent can differentiate into an agent group capable of collaborating to accomplish the target task. In AutoAct, the Meta-Agent can be initialized with any open-source model.
Target Task Information.
In this paper, we mainly focus on agent learning from scratch, which means the task information at hand is quite limited, primarily encompassing three aspects: task name $\mathcal{M}$, task description $\mathcal{P}$, task data examples $\mathcal{C}$. Concretely, $\mathcal{P}$ represents a detailed description of the task’s characteristics, properties, and other relevant information. $\mathcal{C}=\{q_{i},a_{i}\}_{i=1}^{|\mathcal{C}|}$ indicates $|\mathcal{C}|$ question-answer example pairs of the task, where $|\mathcal{C}|$ is very small which users can effortlessly provide (e.g., a few demonstrations). For a more in-depth view of task information, please refer to Appendix B. Note that the task information serves as the only user-provided knowledge of the task for AutoAct to conduct automatic agent learning.
Tool Library.
To facilitate our agents in automatic task planning, we provide a comprehensive tool library at their disposal. The tool library can be denoted as $\mathcal{T}=\{m_{i},d_{i},u_{i}\}_{i=1}^{|\mathcal{T}|}$, where $m$ represents the name of each tool, $d$ defines the functionality of each tool, $u$ details the usage instruction of each tool, and $|\mathcal{T}|$ stands for the tools amount of the library. In our automatic procedure, the Meta-Agent has the autonomy to select appropriate tools from the tool library based on the task information. Users also have the option to expand the tool library according to their specific needs, allowing for more flexible utilization. We list part of our tool library in Appendix C.
2.2 Starting from Scratch via Self-Instruct
To acquire a sufficient amount of task data and provide an ample training resource, it is necessary to augment the data based on the examples at hand. We accomplish this process through self-instruct. Initially, the database $\mathcal{D}$ is set to be equal to the task data examples $\mathcal{C}$, with $\mathcal{C}$ as the seed for data generation. In each round, the Meta-Agent generates new question-answer pairs by few-shot prompting, and the few-shot prompt examples are randomly sampled from $\mathcal{D}$. The generated data will be added to $\mathcal{D}$ followed by filtering, with the exclusion of format erroneous and duplicate data before its inclusion. Eventually, we obtain a database $\mathcal{D}=\{q_{i},a_{i}\}_{i=1}^{|\mathcal{D}|}$, where the number of data $|\mathcal{D}|$ satisfies $|\mathcal{D}|\gg|\mathcal{C}|$. The prompt we use for self-instruct can be seen in Appendix D.1.
2.3 Automatic Agent Learning via Self-Planning
Automatic Tool Selection.
With the tool library at hand, we ask the Meta-Agent to select applicable tools for each task automatically. Specifically, we put $\mathcal{T}=\{m_{i},d_{i},u_{i}\}_{i=1}^{|\mathcal{T}|}$ in the form of a tool list as part of the prompt. Along with $\mathcal{T}$, the prompt also includes the task’s description $\mathcal{C}$. Finally, we instruct the Meta-Agent to select an appropriate set of tools $\mathcal{T}_{s}$ ($\mathcal{T}_{s}\subset\mathcal{T}$) to wait for synthesizing trajectories. The prompt we use for automatic tool selection can be seen in Appendix D.2.
Trajectories Synthesis.
Without relying on closed-source models, we enable the Meta-Agent to synthesize planning trajectories on its own. Equipped with $\mathcal{T}_{s}$, we instruct the Meta-Agent to synthesize trajectories in a zero-shot manner on the database $\mathcal{D}$ adhering to the format of Thought-Action-Observation as defined in Yao et al. (2023). In order to obtain high-quality synthesized trajectories, we filter out all the trajectories with $\texttt{reward}<1$ and collect trajectories with exactly correct answers ($\texttt{reward}=1$) as the training source for self-differentiation. The prompt for trajectories synthesis can be seen in Appendix D.3.
Self-Differentiation.
In order to establish a clear division-of-labor, we leverage synthesized planning trajectories to differentiate the Meta-Agent into three sub-agents with distinct functionalities:
[…]
    Plan-Agent $\pi_{\rm plan}$ undertakes task decomposition and determines which tool to invoke in each planning loop (Eq.2).
[…]
    Tool-Agent $\pi_{\rm tool}$ is responsible for how to invoke the tool (Eq.3) by deciding the parameters for the tool invocation.
[…]
    Reflect-Agent $\pi_{\rm reflect}$ engages in reflection by considering all the historical trajectories and providing a reflection result (Eq.4).
We assume that the planning loop at time $t$ can be denoted as $(\tau_{t},\alpha_{t},o_{t})$, where $\tau$ denotes Thought, $\alpha$ signifies Action, and $o$ represents Observation. $\alpha$ can be further expressed as $(\alpha^{m},\alpha^{p})$, where $\alpha^{m}$ is the name of the action, and $\alpha^{p}$ is the parameters required to perform the action. Then the historical trajectory at time $t$ can be signaled as:
$\mathcal{H}_{t}=(\tau_{0},\alpha_{0},o_{0},\tau_{1},...,\tau_{t-1},\alpha_{t-1},o_{t-1}).$ (1)
Eventually, supposing that the prompts of target task information, planning format requirements, and the question are all combined as $\mathcal{S}$, the responsibilities of each sub-agent can be defined as:
$\tau_{t},\alpha_{t}^{m}$ $=\pi_{\rm plan}(\mathcal{S},\mathcal{T}_{s},\mathcal{H}_{t}),$ (2)
$\alpha_{t}^{p}$ $=\pi_{\rm tool}(\mathcal{S},\mathcal{T}_{s},\mathcal{H}_{t},\tau_{t},\alpha_{t}^{m}),$ (3)
$\tau^{r},\alpha^{r}$ $=\pi_{\rm reflect}(\mathcal{S},\mathcal{T}_{s},\mathcal{H}),$ (4)
where $\tau^{r}$ and $\alpha^{r}$ represent the thought and action of the reflection process respectively, and $\mathcal{H}$ is the planning history after finishing the answer. The trajectories can be reorganized based on the responsibilities above and fed to the Meta-Agent for self-differentiation. Our differentiation is a parameter-efficient fine-tuning process to achieve resource-efficient learning. Particularly, for each sub-agent, we train a specific LoRA Hu et al. (2022).
Group Planning.
After obtaining the task-specific sub-agents, any new question is processed through group planning among the sub-agents to achieve the desired outcome. Once the tool name $\alpha_{t}^{m}$ generated by the Plan-Agent is triggered at time $t$, the Tool-Agent is roused to determine the parameters $\alpha_{t}^{p}$ transferred to the specific tool. The return result of the tool is treated as the observation $o_{t}$ and handed to the Plan-Agent. After the collaboration between the Plan-Agent and Tool-Agent finishes with a prediction, the Reflect-Agent comes on the stage to reflect on the history and provide a reflection result contained in the reflection action $\alpha^{r}$. If the reflection result indicates that the prediction is correct, the whole planning process concludes. Otherwise, the Plan-Agent and Tool-Agent will continue the planning based on the reflection information. The specific sequence of the group planning process can be referred to the trajectory example on the right side of Figure 2.
Backbone Method HotpotQA ScienceQA
Easy Medium Hard All G1-4 G5-8 G9-12 All
[…]
3 Experimental Setup
Tasks.
We evaluate our method on HotpotQA Yang et al. (2018) and ScienceQA Lu et al. (2022). HotpotQA is a multi-hop QA task challenging for rich background knowledge, the answer of which is usually a short entity or yes/no. Following Liu et al. (2023), we randomly select 300 dev questions divided into three levels for evaluation, with 100 questions in each level. For HotpotQA, the $\texttt{reward}\in[0,1]$ is defined as the F1 score grading between the prediction and ground-truth answer. ScienceQA is a multi-modal QA task spanning various scientific topics. We also divide the test set into three levels based on the grade, with 120 randomly sampled data in each level. Since ScienceQA is a multi-choice QA task, the $\texttt{reward}\in\{0,1\}$ is exactly the accuracy. Note that due to the limitations of LMs in generating images, for ScienceQA, during the self-instruct stage, we directly generate captions for the images instead.
Baselines.
We choose the open-source Llama-2 models Touvron et al. (2023) as the backbones of our Meta-Agent and sub-agents. The compared baselines are as follows: 1) CoT Wei et al. (2022), the naive Chain-of-Thought reasoning method. 2) ReAct Yao et al. (2023), a well-known single-agent framework based on few-shot learning that performs planning and action iteratively. 3) Chameleon Lu et al. (2023), another few-shot single-agent framework that performs planning before action. 4) Reflexion Shinn et al. (2023), a single-agent framework to reinforce language agents through linguistic feedback. 5) BOLAA Liu et al. (2023), a multi-agent framework that customizes different agents through prompts. 6) FireAct Chen et al. (2023a), a single-agent framework with fine-tuning on diverse kinds of trajectories generated by GPT-4 OpenAI (2023). 7) GPT-3.5-Turbo OpenAI (2022). To ensure fairness, we maintain an equal training trajectory volume of 200 for FireAct and AutoAct (200 synthesized data). As Reflexion provides answer correctness labels during reflection but other methods including AutoAct do not, we test all the other methods twice and choose the correct one for evaluation. For all the prompt-based baselines, we uniformly provide two examples in the prompt.
Training Setups.
We fine-tune all our models with LoRA Hu et al. (2022) in the format proposed in Alpaca Taori et al. (2023). Our fine-tuning framework leverages FastChat Zheng et al. (2023) using DeepSpeed Rasley et al. (2020). We set the learning rate of 1e-4 and the sequence length of 4096 for all the model scales. The training epoch for the 13b and 7b models is 5 and for the 70b model is 3. The batch size is 4 for the 13b and 7b models and 1 for the 70b model. We use the AdamW optimizer Loshchilov and Hutter (2019) with a cosine learning scheduler. We detail the hyper-parameters for training in Appendix A.
4 Results
Compare to Prompt-based Agent Learning Baselines.
As shown in Table 2, the 13b and 70b models consistently outperform various prompt-based baselines. The 70b model even surpasses the agent performance of GPT-3.5-Turbo, achieving a rise of $\uparrow$3.77on HotpotQA and $\uparrow$6.39on ScienceQA. The performance of the 7b model is comparable to other methods to some extent. However, it fails to deliver the same impressive results due to its limited self-instruct and self-planning capabilities. Therefore, whether in a single-agent or multi-agent architecture, prompt-based methods relying on few-shot demonstrations fail to precisely customize the behavior of the agent, which is also supported by the fact that FireAct widely outperforms ReAct and BOLAA in the context of iterative planning. In addition, our investigation reveals a visible disparity in open-source models between the performance of many prompt-based planning baselines (relying on various external tools) and CoT (relying on the models’ intrinsic reasoning abilities). This discrepancy underscores the formidable challenge of unlocking planning capabilities by prompting.
Compare to Fine-tuning-based Agent Learning Baselines.
Further focusing on FireAct in Table 2, despite the assistance of GPT-4, FireAct’s approach of assigning the entire planning task to a single model proves to be burdensome. As a result, its performance on ScienceQA even falls short compared to the prompt-based global planning method, Chameleon. AutoAct employs self-differentiation to decouple the planning process and reaches a clear division-of-labor among sub-agents for group planning, resulting in a improvement than FireAct, with an enhancement of $\uparrow$5.77on HotpotQA and $\uparrow$6.67on ScienceQA with 70b model. Additionally, AutoAct achieves self-planning without relying on closed-source models and large-scale labeled datasets, which paves the way for automatic agent learning with open-source models from scratch. In ablation study (§4) and human evaluation (§5), we will further validate that the quality of trajectories synthesized by AutoAct is not inferior to FireAct trained on trajectories synthesized using GPT-4.
Figure 3: Performance of AutoAct on different training data scales. (a-c) represents the results of the model trained on self-synthesized trajectories. (d-f) represents the results of the model trained on trajectories synthesized by a stronger model, where the dashed line is the baseline trained on self-synthesized trajectories.
Figure 4: Performance of AutoAct based on different degrees of labor division. One is training a single model with all the differentiated data. Three represents the differentiation into three agents: plan, tool, and reflect. Tool Specified indicates further differentiating the tool-agent with one tool, one agent.
HotpotQA ScienceQA
  AutoAct 48.47 78.61
[…]
Table 3: Approach ablations of AutoAct. - reflection symbolizes removing the reflect-agent in AutoAct. - multi denotes feeding all the differentiated data into one model for fine-tuning. - fine-tuning indicates zero-shot prompt planning with the three agents defined in AutoAct. - filtering represents self-differentiation on all the trajectories generated in zero-shot planning without filtering wrong cases.
Single-agent Learning vs. Multi-agent Learning.
Under identical settings, multi-agent architectures generally exhibit better performance than single-agent (ReAct vs. BOLAA, FireAct vs. AutoAct), which aligns with Simon’s theory of bounded rationality. Seemingly contrary to expectations, despite being a single-agent architecture, Chameleon outperforms BOLAA (even FireAct on ScienceQA). However, we analyze that this can be attributed to the way it leverages tools. In Chameleon, the process of deciding tool parameters is considered a form of tool invocation, and specialized few-shot prompts are designed to guide the model through this process. From this aspect, Chameleon, despite being nominally a single-agent architecture, exhibits characteristics that resemble a multi-agent architecture, which does not contradict our initial conclusion. Indeed, we can also explain from the perspective of optimizing objectives. Another well-known economic principle, Goodhart’s Law Goodhart (1984), states that “When a measure becomes a target, it ceases to be a good measure”. This implies that optimizing one objective on the same agent will inevitably harm other optimization objectives to some extent. Therefore, it is not optimal to optimize all objectives on a single agent, and a multi-agent architecture happens to address this issue. However, we analyze in §5 that excessive fine-grained division-of-labor is not the best approach and a moderate division of labor benefits group performance.
Approach Ablations.
Table 3 presents the performance of AutoAct on the 70b model after removing certain key processes. It can be observed that the least impactful removal is the - reflect. We investigate that in the zero-shot scenario, the model tends to be over-confident in its answers. It typically only recognizes its errors when there are obvious formatting mistakes or significant repetitions in the planning process. Consistent with previous findings, the removal of the - multi agents leads to a noticeable decrease in performance. A more exciting discovery is that the results of - multi are comparable to those of FireAct. This indirectly suggests that the trajectory quality generated by the 70b model may be no worse than that of GPT-4. As expected, the performance deteriorates after - fine-tuning, which once again confirms the previous conclusion. To demonstrate the necessity of filtering out planning error data, we specifically remove the filtering process (- filtering) to examine the performance of AutoAct. The results indicate that the damage caused by training on unfiltered data is even greater than that of - fine-tuning.
Figure 5: Case study. AutoAct (b) successfully addresses the failure in ReAct (a) by employing a more scientific combination of tools and making more accurate tool invocations. With more planning rounds, AutoAct (c) can validate its inner answers by continuing more rounds of self-verification. While this can also lead to a longer context, gradually deviating AutoAct (d) from the original question.
Figure 6: Human evaluation of trajectories generated by Llama-2-70b-chat on HotpotQA. We compare the number of planning rounds, the logical correctness of thoughts, action types, action parameters, and the overall coherence of each trajectory.
5 Analysis
Larger training data scale does not necessarily mean better results.
We evaluate the influence of different training data scales on the performance of self-planning in Figure 3(a-c). It can be observed that the overall performance of different models steadily improves as the training data scale increases. However, it goes to stability with minimal fluctuations once the data scale exceeds 200. We speculate that this may be due to the limited ability of naive self-instruct to boost internal knowledge of the language model. The process from self-instruct to self-planning can be seen as the distillation and rumination of internal model knowledge. As the training data increases, the knowledge which can be extracted through self-instruct decreases. Despite our efforts to filter out duplicate data, similar data will inevitably reoccur. So the mindless increase in data scale can lead to a significant surge in similar data, which undermines the benefits of increasing the data scale and ultimately makes it challenging to improve model performance or even leads to over-fitting (see the slight decrease in overall performance on 70b at 250 and 300). To further confirm the role of training data, we decouple the models from the training data and evaluate their training results on trajectories synthesized by other stronger models. From Figure 3(d-f), we can see consistent conclusions with previous findings. The performance improvement becomes increasingly challenging beyond a dataset size of 200, regardless of the size matching between the backbone model and the data-synthetic model. Therefore, maximizing the diversity of the synthesized data in the database may be a key improvement direction for AutoAct. Some previous works Xu et al. (2023); Yu et al. (2023); Li et al. (2023) have attempted to improve upon the naive self-instruct, but none of them have focused on better mobilizing the language model’s internal knowledge without external information, and we leave this for our future work.
Moderate division of labor benefits group planning performance.
To explore the impact of the different granularity of self-differentiation and group planning, we further subdivide the tool agent, assigning dedicated agents to manipulate each specific tool. We contrast the performance of One agent, Three agents (vanilla AutoAct), and the Tool-Specified setting in Figure 4. It can be observed that finer task allocation does not necessarily lead to better performance. This is consistent with the findings in Qiao et al. (2023a) which indicate that multi-tool joint learning often outperforms single-tool individual learning. Therefore, appropriate differentiation (Three) can alleviate the pressure on individual agents, aligning with Simon’s principle of bounded rationality. However, excessive differentiation (Tool-Specified) not only fails to achieve better results but can sometimes even be less effective than not differentiating (One) at all. Moreover, it appears that the performance loss of tool-specific agents compared to the three-agent approach is more significant on harder problems. This is because challenging problems typically require more planning steps and higher levels of collaboration among tools. By unifying tool invocations under one agent, it becomes possible to effectively learn the interplay and interconnectedness between tools, thereby compensating for potential information gaps arising from using tool-specific agents.
Human Evaluation.
To get a deeper understanding of the capability of AutoAct, we manually compare the quality of trajectories generated by different methods from five aspects. We ask five NLP experts to individually select the optimal trajectories generated by all methods in terms of the number of planning rounds, the logical correctness of thoughts, action types, action parameters, and overall coherence. The final results are determined based on major votes. During the evaluation, it is hidden for the evaluators of the correspondence between the trajectories and the methods. We delete the reflection-related parts from the trajectories generated by AutoAct and randomly shuffle the order of trajectories of each method in each data to minimize the potential bias as much as possible.
The evaluation results are depicted in Figure 6 and we further provide some cases in Figure 5. We can observe a clear advantage for AutoAct over other methods in determining the action type and action parameters. This indicates that decoupling the missions of planning and tool invocation can lead to better performance for both, alleviating the overwhelming pressure on a single agent. A more intuitive comparison can be observed in Figure 5 (a)(b). AutoAct successfully addresses the failure in ReAct by employing a more scientific combination of tools and making more accurate tool invocations. Furthermore, AutoAct tends to consume more planning rounds than other methods. This allows AutoAct to perform better on harder problems. However, this characteristic can be a double-edged sword when it comes to simple problems. A surprising aspect is that AutoAct can validate its inner (thought) answers by continuing more rounds of self-verification (see Figure 5 (c)). Unfortunately, this can also lead to a longer context, gradually deviating AutoAct from the original question (see Figure 5 (d)).
6 Related Work
LLM-Powered Agents.
The rise of LLMs has positioned them as the most promising key to unlocking the door to Artificial General Intelligence (AGI), providing robust support for the development of LLM-centered AI agents Wang et al. (2023b); Xi et al. (2023); Wang et al. (2023d, e). Related works focus primarily on agent planning Yao et al. (2023); Song et al. (2022); Chen et al. (2023a), external tools harnessing Patil et al. (2023); Qiao et al. (2023a); Qin et al. (2023), collective intelligence among multi-agents Liang et al. (2023); Liu et al. (2023); Chen et al. (2023b), human and social property inside agents Zhang et al. (2023a); Park et al. (2023); Mao et al. (2023), etc. However, despite their success, existing methods still face two major troubles. Firstly, most agents heavily rely on prompts for customization, which makes it difficult to precisely tailor the behavior of the agent, resulting in unexpected performance at times. Secondly, each agent is compelled to master all skills, making it challenging for the agent to achieve expertise in every domain. In response, our approach leverages a proper division-of-labor strategy and fine-tuning each sub-agent to equip different agents with distinct duties. These agents collaborate to accomplish tasks orderly and effectively.
Agent Fine-Tuning.
Despite the vast interest in LLM-powered agents, the construction of agents through fine-tuning has received limited attention. Most early works concentrate on fine-tuning to optimize the model’s reasoning capabilities Liu et al. (2022); Fu et al. (2023) or tool proficiency Patil et al. (2023); Qiao et al. (2023a); Qin et al. (2023). Recently, Chen et al. (2023a) attempt to fine-tune agents with diverse tasks and trajectories for a better planning capability. Zeng et al. (2023) apply a hybrid instruct-tuning strategy that enhances the agent’s abilities while preserving its generalization. However, these methods still require a model to be a generalist. Moreover, the trajectories in the training data are annotations from GPT-3.5/GPT-4 OpenAI (2022, 2023), which incurs significant costs. Our approach enables the Meta-Agent to autonomously synthesize trajectories and achieve self-planning in a zero-shot manner, without relying on closed-source models or human labor.
7 Conclusion
In this paper, we propose AutoAct, an automatic agent learning framework that does not rely on large-scale annotated data and synthetic trajectories from closed-source models, while alleviating the pressure on individual agents by explicitly dividing the workload, thereby enhancing the collective performance of multi-agents. Experimental results demonstrate that AutoAct performs superior on challenging question-answering benchmarks compared to various strong baselines, and AutoAct with Llama-2-13b model can even obtain comparable performance of zero-shot GPT-3.5-Turbo agent. Extensive analysis indicates the effectiveness of our appropriate division-of-labor strategy, with the trajectory quality generated by AutoAct significantly outperforming that of other methods from multiple aspects.
Limitations
In this paper, we focus on constructing an automatic agent learning framework dubbed AutoAct. Despite our best efforts, this paper may still have some remaining limitations.
Tasks and Backbones.
For experimental convenience, we only evaluate our method on complex question-answering tasks with the Llama-2-chat model series. However, there are many other interactive scenarios and backbone models beyond these. Other complex tasks include web Yao et al. (2022); Zhou et al. (2023b), household Puig et al. (2018); Shridhar et al. (2021), robotics Ichter et al. (2022), etc., and more backbone models include Vicuna Zheng et al. (2023), ChatGLM Du et al. (2022), Mistrial Jiang et al. (2023), etc. We plan to conduct further research on applying AutoAct to a wider range of tasks and backbones in the future.
Boosting Knowledge via Self-Instruct.
As analyzed in §5, the planning performance of AutoAct can be limited by the model’s ability to access internal knowledge through self-instruct. While the current phenomenon allows us to achieve lightweight self-differentiation in terms of parameters and data, it is still necessary to research how to enrich knowledge as much as possible within the constraints of limited data.
Self-Improvement.
Recent research has shed light on self-improvement techniques that enhance LLMs by iteratively training them on self-synthesized data Zelikman et al. (2022); Huang et al. (2023); Gülçehre et al. (2023); Aksitov et al. (2023). This approach allows the model to continually learn and refine its performance on its own. Our approach also involves training on self-synthesized data and we believe that further using the iterative thinking of self-improvement will significantly enhance the performance of our method.
Appendix A Hyper-Parameters
Name Llama-2-7b&13b-chat Llama-2-70b-chat
  lora_r 8 8
lora_alpha 16 16
lora_dropout 0.05 0.05
lora_target_modules q_proj, v_proj q_proj, v_proj
model_max_length 4096 4096
per_device_batch_size 2 2
gradient_accumulation_steps 1 1
warmup_ratio 0.03 0.03
epochs 5 3
batch size 4 1
learning rate 1e-4 1e-4
Table 4: Detailed hyper-parameters we use for training.
See Table 4.
Appendix B Task Information
Task Name: HotpotQA
Task Description
: This is a question-answering task that includes high-quality multi-hop questions. It tests language modeling abilities for multi-step reasoning and covers a wide range of topics. Some questions are challenging, while others are easier, requiring multiple steps of reasoning to arrive at the final answer.
Task Data Examples:
Question: From 1969 to 1979, Arno Schmidt was the executive chef of a hotel located in which neighborhood in New York?
Answer: Manhattan
Question: Are both Shangri-La City and Ma’anshan cities in China?
Answer: yes
Task Name: ScienceQA
Task Description
: This is a multimodal question-answering task that necessitates a model to utilize tools for transforming image information into textual data. Simultaneously, this task incorporates substantial background knowledge, requiring the language model to acquire external information to enhance its comprehension of the task.
Task Data Examples:
Question: Which of these states is the farthest north?
Options: (A) West Virginia (B) Louisiana (C) Arizona (D) Oklahoma
Caption: An aerial view of a painting of a forest.
Answer: A. West Virginia
Question: Identify the question that Tom and Justin’s experiment can best answer.
Context: The passage below describes an experiment. Read the passage and then follow the instructions below. Tom placed a ping pong ball in a catapult, pulled the catapult’s arm back to a 45 angle, and launched the ball. Then, Tom launched another ping pong ball, this time pulling the catapult’s arm back to a 30 angle. With each launch, his friend Justin measured the distance between the catapult and the place where the ball hit the ground. Tom and Justin repeated the launches with ping pong balls in four more identical catapults. They compared the distances the balls traveled when launched from a 45 angle to the distances the balls traveled when launched from a 30 angle. Figure: a catapult for launching ping pong balls.
Options: (A) Do ping pong balls stop rolling along the ground sooner after being launched from a 30-angle or a 45-angle? (B) Do ping pong balls travel farther when launched from a 30-angle compared to a 45-angle?
Caption: A wooden board with a wooden head on top of it.
Answer: B. Do ping pong balls travel farther when launched from a 30 angle compared to a 45 angle?
Appendix C Tool Library
See Table 5.
Name Definition Usage
  BingSearch
BingSearch engine can search for rich knowledge on the internet based on keywords, which can compensate for knowledge fallacy and knowledge outdated.
BingSearch[query], which searches the exact detailed query on the Internet and returns the relevant information to the query. Be specific and precise with your query to increase the chances of getting relevant results. For example, Bingsearch[popular dog breeds in the United States]
Retrieve
Retrieve additional background knowledge crucial for tackling complex problems. It is especially beneficial for specialized domains like science and mathematics, providing context for the task
Retrieve[entity], which retrieves the exact entity on Wikipedia and returns the first paragraph if it exists. If not, it will return some similar entities to retrieve. For example, Retrieve[Milhouse]
Lookup
A Lookup Tool returns the next sentence containing the target string in the page from the search tool, simulating Ctrl+F functionality on the browser.
Lookup[keyword], which returns the next sentence containing the keyword in the last passage successfully found by Retrieve or BingSearch. For example, Lookup[river].
Image2Text
Image2Text is used to detect words in images convert them into text by OCR and generate captions for images. It is particularly valuable when understanding an image semantically, like identifying objects and interactions in a scene.
Image2Text[image], which generates captions for the image and detects words in the image. You are recommended to use it first to get more information about the image to the question. If the question contains an image, it will return the caption and OCR text, else, it will return None. For example, Image2Text[image].
Text2Image
Text2Image Specializes in converting textual information into visual representations, facilitating the incorporation of textual data into image-based formats within the task.
Text2Image[text], which generates an image for the text provided by using multimodal models. For example, Text2Image[blue sky]
Code Interpreter
Code Interpreter is a tool or software that interprets and executes code written in Python. It analyzes the source code line by line and translates it into machine-readable instructions or directly executes the code and returns Execution results
Code[python], which interprets and executes Python code, providing a line-by-line analysis of the source code and translating it into machine-readable instructions. For instance, Code[print("hello world!")]
Table 5: Part of our tool library.
Appendix D Prompt
D.1 Prompt for Self-Instruct
See Table 6.
Prompt for Self-Instruct
I want you to be a QA pair generator to generate high-quality questions for use in Task described as follows:
Task Name: [task_name]
Task Description: [task_description]
Here are some Q&A pair examples from the Task:
[QA_pairs]
Modeled on all the information and examples above, I want you to generate new different [gen_num_per_round] Question-Answer pairs that cover a wide range of topics, some of which are difficult, some of which are easy, and require multiple steps of reasoning to get to the final answer. The format is like below:
[one_example]
Table 6: Prompt used for self-instruct.
D.2 Prompt for Tool Selection
See Table 7.
Prompt for Automatic Tool Selection
To successfully complete a complex task, the collaborative effort of three types of agents is typically required:
1. Plan Agent. This agent is used to plan the specific execution process of the benchmark, solving a given task by determining the order in which other expert language models are invoked;
2. Tool Agent. This agent is employed to decide how to use a specific tool when addressing a task. Tools encompass interactive tools within the task environment as well as external tools or models. The Tool Agent includes various tools that can be flexibly chosen;
3. Reflect Agent. This agent reflects on historical information and answers to assess whether the response aligns with the provided query.
Above all, the Tool Agent includes many tools that can be flexibly selected. Now your task is to select 3 tools from the Tool Library for solving a given task. Note that all tools are based on language models, and their inputs and outputs must be text. You only need to provide the names and descriptions of the tools in order, without any additional output.
Task Prompt Template
The following is the given task name and description, and you need to choose 3 corresponding tools from the Tool Library according to the above rules in the format of one line, one tool.
Task Name: [task_name]
Task Description: [task_description]
Tool Library: [list_of_tools]
Table 7: Prompt used for automatic tool selection.
D.3 Prompt for Trajectories Synthesis
See Table 8.
Prompt for Trajectories Synthesis
I expect you to excel as a proficient question answerer in the task.
Task Name: [task_name]
Task Description: [task_description]
Solve a question-answering task with interleaving Thought, Action, and Observation steps. Thought can reason about the current situation, and Action can be [action_num] types:
list of action selected from automatic tool selection [name, definition , usage]
Question: [question][scratchpad]
Table 8: Prompt used for trajectories synthesis.